eviant:

Find the slope of the line that passes through the pairs of points. (2,6) (7,0)

eviant:

@563blackghost

Ultrilliam:

@Shadow

563blackghost:

\(\Large\bf{\frac{y_{2}-y_{1}}{x_{2}-x_{1}} \rightarrow \frac{0-6}{7-2}}\)

eviant:

So, \[\frac{ -6 }{ 5 }\]

563blackghost:

yup, but you would write it as \(\large\bf{-\frac{6}{5}}\).

eviant:

ok, thanks!